# Secret-Key Generation from Private Identifiers under Channel Uncertainty

### v1

@@ A-A Proof of Proposition 1
where (a) follows from (36), shown below, and the codebook $\mathcal{C}_{n}$ is independent of $(\tilde{\mathbf{X}}^{n},X^{n},\mathbf{Y}^{n}_{k},\mathbf{Z}^{n}_{l})$, (b) holds because $V^{n}$ is a function of $(J_{v_{1}},J_{v_{2}})$, and (c) follows because $H(Z^{n}_{l}|V^{n},\mathcal{C}_{n})\leq n(H(Z_{l}|V)+\gamma^{\prime}_{n})$ and $\epsilon^{\prime\prime}_{n}~{}\triangleq~{}\gamma^{\prime}_{n}+\epsilon^{\prime}_{n}$. For brevity, define $E~{}\triangleq~{}(J_{u_{2}},J_{u_{3}},J_{v_{2}})$, where the decoder can reliably estimate the index $E$ for given $(\mathbf{Y}^{n}_{k},J_{u_{1}},J_{v_{1}})$. Observe that
@@
where (a) follows since $U^{n}$ is a function of $(J_{u_{1}},J_{u_{2}},J_{u_{3}},J_{v_{1}},J_{v_{2}})$, (b) is due to the Markov chain $E-(X^{n},J_{u_{1}},J_{v_{1}})-\mathbf{Y}^{n}_{k}$ and conditioning reduces entropy, (c) follows from Fano’s inequality with $\epsilon_{n}\downarrow 0$ as $\delta\downarrow 0$ and $n\rightarrow\infty$, and (d) because $H(\tilde{\mathbf{X}}^{n}|X^{n},U^{n},\mathcal{C}_{n})\leq n(H(\tilde{\mathbf{X}}|X,U)+\gamma^{\prime\prime}_{n})$ for jointly typical sequences (cf. [44, eq. (16)]) and $\epsilon^{\prime}_{n}~{}{\triangleq}~{}\gamma^{\prime\prime}_{n}+\epsilon_{n}$.
@@ A-B Proof of Proposition 2
The cardinality bounds of the auxiliary random variables can be obtained from the support lemma [43, Lemma 3.4].
We show the proof for the GS model via a result derived in [20] in the absence of action cost. By replacing (6) with $I(X^{n};J,\mathbf{Z}^{n}_{l})\leq n(R_{L}+\delta)$, the outer bound for a pair $(k,l)$, denoted by $\mathcal{O}_{G_{kl}}$, is [20, Th. 3]
@@
In Definition 3, the constraints on the secret-key and storage rates are the same as in [20, Def. 6], and the resulting bounds are given in the same form as in (37). For the privacy-leakage rate, we expand the right-hand side of $R_{L}$ in (37) as
@@
where we used the Markov chain $V-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$. By (38), the privacy-leakage rate is lower bounded as
@@
$\geq n(I(X;U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};V)-I(\mathbf{Z}_{l};V)),$ (39)
@@
$\mathcal{R}_{G}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\mathcal{O}_{G_{kl}}.$ (41)
The proof for the CS model can be derived using the same reasoning using [20, Th. 4] instead of [20, Th. 3].
@@ Appendix B Proof of Lemma 1
Note that (2) depends on the marginal distribution of $(X,\mathbf{Y}_{k})$, (3) depends on the marginal distribution of $X$, and (5) and (6) depend on the marginal distribution of $(X,\mathbf{Z}_{l})$. Therefore, without loss of generality, using [45, Th. 3.5.2] and (42), it suffices to consider
@@ C-A Achievability Proof
Lemma 4.
(Weinstein–Aronszajn Identity [47, Appx. B]) For any $a\in\mathbb{R}_{+}$ and matrix $\mathbf{H}\in\mathbb{R}^{\Omega\times 1}$, we have
${\det}(\mathbf{H}a\mathbf{H}^{\intercal}+{\bf I}_{\Omega})=a\mathbf{H}^{\intercal}\mathbf{H}+1,$
@@
$f(\mathbf{H}^{\intercal}\mathbf{H})=\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}\mathbf{H}+1}{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}\mathbf{H}+1}\right)$
@@
$\mathbf{Y}_{k}$ $=\mathbf{H}_{k}U+\mathbf{H}_{k}\Theta+\mathbf{N}_{\mathbf{Y}_{k}},$
$\mathbf{Z}_{l}$ $=\tilde{\mathbf{H}}_{l}U+\tilde{\mathbf{H}}_{l}\Theta+\mathbf{N}_{\mathbf{Z}_{l}}.$
@@
$I(\mathbf{Y}_{k};U)$ $=\frac{1}{2}\log\frac{\sigma^{2}_{X}\mathbf{H}_{k}^{\intercal}\mathbf{H}_{k}+1}{\alpha\sigma^{2}_{X}\mathbf{H}_{k}^{\intercal}\mathbf{H}_{k}+1},$
$I(\mathbf{Z}_{l};U)$ $=\frac{1}{2}\log\frac{\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1}{\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1}$
@@ C-B Converse Proof
We will need the following lemmas. These lemmas convert vector observations in (16) into scalar Gaussian random variables using sufficient statistics [39, Sect. 2.9]. This transformation plays an important role in deriving the outer bound of Gaussian sources.
Lemma 6.
([48, Lemma 3.1]) Consider a channel with input $W$ and output $\tilde{\bf W}$, namely, $\tilde{\bf W}~{}{\triangleq}~{}{\bf A}W+{\bf N}_{\tilde{\bf W}},$ where ${\bf N}_{\tilde{\bf W}}\sim\mathcal{N}({\bf 0},{\bf\Sigma}_{\tilde{\bf W}})$. A sufficient statistic to correctly determine $W$ from $\tilde{\bf W}$ is the following scalar
@@
As $(X,\bar{Y}_{k},\bar{Z}_{l})$ are scalar Gaussian random variables, when the squared value of the correlation coefficient of $(X,\bar{Y}_{k^{*}})$ is greater than that of $(X,\bar{Z}_{l^{*}})$, i.e., $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, implying that $\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}\geq\tilde{\mathbf{H}}^{\intercal}_{l}\mathbf{H}_{l}$ for any pair $(k,l)$, there exist Gaussian random variables $(X^{\prime},\bar{Y}^{\prime}_{k},\bar{Z}^{\prime}_{l})$ such that $X^{\prime}-\bar{Y}^{\prime}_{k}-\bar{Z}^{\prime}_{l}$ is satisfied, and the marginal distributions of $(X,\bar{Y}_{k})$ and $(X^{\prime},\bar{Y}^{\prime}_{k})$ and that of $(X,\bar{Z}_{l})$ and $(X^{\prime},\bar{Z}^{\prime}_{l})$ coincide [40, Lemma 6]. In the subsequent discussions, we denote $(X^{\prime},\bar{Y}^{\prime}_{k},\bar{Z}^{\prime}_{l})$ as $(X,\bar{Y}_{k},\bar{Z}_{l})$ for brevity. This property is used in the derivation of the following lemma.
@@
If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}$, the outer bounds of the GS and CS models are provided as
$\mathcal{R}_{G}~{}$ $\subseteq~{}\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bar{\mathcal{O}}_{G_{kl}},~{}~{}~{}\mathcal{R}_{C}\subseteq~{}\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bar{\mathcal{O}}_{C_{kl}},$ (57)
@@
and $U$ satisfies the Markov chain $U-X-\bar{Y}_{k}-\bar{Z}_{l}$. If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}<\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}$, $\mathcal{R}_{G}=\mathcal{R}_{C}=\{(R_{S},R_{J},R_{L}):R_{S}=0,R_{J}\geq 0,R_{L}\geq 0\}.$
@@
See Appendix D. ∎
@@
The conditional Fisher information of $A$ is defined by $\mathbb{J}(A|U)=\mathbb{E}\left[\left(\frac{\partial\log f_{A|U}(a|u)}{\partial a}\right)^{2}\right],$ where the expectation is taken over $(U,A)$ [42, Def. 1].
@@
([42, Cor. 1]) Let $W,A,B$ be random variables, and let the density for any combination of them exist. Moreover, assume that given $W$, $A$ and $B$ are independent. Then, we have
Lemma 10.
For $0\leq r\leq 1/\bar{\nu}_{\bar{Y}_{k}}$, it holds that
@@
From [42, Lemma 3], it follows that
[…]
@@
[…]
=$\overset{\rm(a)}{=}\frac{1}{2}\int_{1/\nu_{\bar{Y}_{k}}}^{1/\nu_{\bar{Z}_{l}}}\mathbb{J}(X+\sqrt{r}N+\sqrt{t-r}N^{\prime}|U)dt$
@@
[…]
@@
where (a) follows by picking a real number $r$ in the range of $0\leq r\leq 1/\bar{\nu}_{\bar{Y}_{k}}$ and using independent Gaussian random variables $N\sim\mathcal{N}(0,1)$ and $N^{\prime}\sim\mathcal{N}(0,1)$, and (b) is due to Lemma 9. Lastly, comparing (64) and (68), we obtain (67). ∎
@@
$\mathcal{R}_{G}$ $\subseteq\bigcup_{0<\alpha\leq 1}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
@@ Appendix D Proof of Lemma 8
where the left-hand sides of (73) and (74), and (75) hold because $\bar{Y}^{n}_{k}$ and $\bar{Z}^{n}_{l}$ are functions of $\mathbf{Y}^{n}_{k}$ and $\mathbf{Z}^{n}_{l}$, respectively, by Lemma 7, and the right-hand sides of (73) and (74) are due to the sufficient statistic [39, Sect. 2.9]. In addition, for the scalar random variables $(X,\bar{Y}_{k},\bar{Z}_{l})$, the Markov chain $X^{n}-\bar{Y}^{n}_{k}-\bar{Z}^{n}_{l}$ holds for any pair $(k,l)$ when $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$. Combining this with (75) gives
@@
=$\overset{\rm(b)}{=}H(S|J,\mathbf{Z}^{n}_{l},\bar{Z}^{n}_{l})-H(S|J,\mathbf{Y}^{n}_{k},\bar{Y}^{n}_{k})+n(\delta+\delta_{n})$
=$\overset{\rm(c)}{=}H(S|J,\bar{Z}^{n}_{l})-H(S|J,\bar{Y}^{n}_{k})+n(\delta+\delta_{n})$
@@
=$\overset{\rm(d)}{=}\sum_{t=1}^{n}\{I(\bar{Y}_{k,t};U_{t}|V_{t})-I(\bar{Z}_{l,t};U_{t}|V_{t})\}+n(\delta+\delta_{n})$ (80)
=$\overset{\rm(e)}{=}\sum_{t=1}^{n}\{I(\bar{Y}_{k,t};U_{t})-I(\bar{Z}_{l,t};U_{t})$
@@
where (a) is due to (5) and Fano’s inequality with $\delta_{n}$ defined in (79) as the secret key $S$ can be reliably estimated from $(J,\mathbf{Y}^{n}_{k})$, (b) follows from the left-hand sides of (73) and (74), (c) holds by the right-hand sides of (73) and (74), (d) follows by [49, Lemma 4.1], (e) follows from the Markov chains $V_{t}-U_{t}-\bar{Y}_{k,t}$ and $V_{t}-U_{t}-\bar{Z}_{l,t}$, and (f) is due to (78), which results in $I(\bar{Y}_{k,t};V_{t})-I(\bar{Z}_{l,t};V_{t})\geq 0$.
@@
=$\overset{\rm(c)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k})-n\delta_{n}$
=$\overset{\rm(d)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k})-n\delta_{n}$ (82)
=$\overset{\rm(f)}{=}\sum_{t=1}^{n}\{h(X_{t}|\bar{Y}_{k,t})-h(X_{t}|J,S,X^{t-1},\bar{Y}^{n}_{k},\bar{Z}^{t-1}_{l})\}-n\delta_{n}$
@@
$I(X^{n};J|\mathbf{Z}^{n}_{l})$ =$\overset{\rm(a)}{=}I(X^{n};J)-I(\mathbf{Z}^{n}_{l};J)$
=$\overset{\rm(b)}{=}I(X^{n};J)-I(\mathbf{Z}^{n}_{l},\bar{Z}^{n}_{l};J)$
=$\overset{\rm(c)}{=}I(X^{n};J)-I(\bar{Z}^{n}_{l};J)$
=$\overset{\rm(d)}{=}I(X^{n};J|\bar{Z}^{n}_{l}),$ (84)
@@
=$\overset{\rm(b)}{=}I(X^{n};J,S,\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-n\delta_{n}$
=$\overset{\rm(c)}{=}I(X^{n};J,S,\bar{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\bar{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-n\delta_{n}$
=$\overset{\rm(d)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k},\bar{Z}^{n}_{l})+I(X^{n};\bar{Y}^{n}_{k}|\bar{Z}^{n}_{l})$
@@
where (a) is due to (84), (b) follows from (54), i.e., $\bar{Y}^{n}_{k}$ is a function of $\mathbf{Y}^{n}_{k}$, (c) holds because the Markov chain $X^{n}-(J,S,\bar{Y}^{n}_{k},\bar{Z}^{n}_{l})-\mathbf{Y}^{n}_{k}$ holds, from the right-hand side of (73), (d) and (e) are due to the Markov chain $(J,S)-X^{n}-Y^{n}_{k}-Z^{n}_{l}$, obtained from (76), and (e) follows by the same steps from (82) to (83).

### v2

@@ A-A Proof of Proposition 1
where (a) follows from (46), shown below, and the codebook $\mathcal{C}_{n}$ is independent of $(\tilde{\mathbf{X}}^{n},X^{n},\mathbf{Y}^{n}_{k},\mathbf{Z}^{n}_{l})$, (b) holds because $V^{n}$ is a function of $(J_{v_{1}},J_{v_{2}})$, and (c) follows because $H(\mathbf{Z}^{n}_{l}|V^{n},\mathcal{C}_{n})\leq n(H(\mathbf{Z}_{l}|V)+\gamma^{\prime}_{n})$ and $\epsilon^{\prime\prime}_{n}~\triangleq~\gamma^{\prime}_{n}+\epsilon^{\prime}_{n}$. For brevity, define $E~\triangleq~(J_{u_{2}},J_{u_{3}},J_{v_{2}})$, where the decoder can reliably estimate the index $E$ for given $(\mathbf{Y}^{n}_{k},J_{u_{1}},J_{v_{1}})$. Observe that
@@
where (a) follows since $U^{n}$ is a function of $(J_{u_{1}},J_{u_{2}},J_{u_{3}},J_{v_{1}},J_{v_{2}})$, (b) is due to the Markov chain $E-(X^{n},J_{u_{1}},J_{v_{1}})-\mathbf{Y}^{n}_{k}$ and conditioning reduces entropy, (c) follows from Fano’s inequality with $\epsilon_{n}\downarrow 0$ as $\delta\downarrow 0$ and $n\rightarrow\infty$, and (d) because $H(\tilde{\mathbf{X}}^{n}|X^{n},U^{n},\mathcal{C}_{n})\leq n(H(\tilde{\mathbf{X}}|X,U)+\gamma^{\prime\prime}_{n})$ for jointly typical sequences (cf. [53, eq. (16)]) and $\epsilon^{\prime}_{n}~{\triangleq}~\gamma^{\prime\prime}_{n}+\epsilon_{n}$.
@@ A-B Proof of Proposition 2
The cardinality bounds of the auxiliary random variables can be obtained from the support lemma [51, Lemma 3.4].
We show the proof for the GS model via a result derived in [21] in the absence of action cost. By replacing (6) with $I(X^{n};J,\mathbf{Z}^{n}_{l})\leq n(R_{L}+\delta)$, the outer bound for a pair $(k,l)$, denoted by $\mathcal{O}_{G_{kl}}$, is [21, Th. 3]
@@
In Definition 3, the constraints on the secret-key and storage rates are the same as in [21, Def. 6], and the resulting bounds are given in the same form as in (47). For the privacy-leakage rate, we expand the right-hand side of $R_{L}$ in (47) as
@@
where we use the Markov chain $V-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$. By (48), the privacy-leakage rate is lower bounded as
@@
$\geq n(I(X;U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};V)-I(\mathbf{Z}_{l};V)).$ (49)
@@
$\mathcal{O}_{G_{kl}}$$\triangleq$} $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$~
@@
$\mathcal{R}_{\rm G}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\mathcal{O}_{G_{kl}}.$ (51)
The proof for the CS model can be derived using the same reasoning from [21, Th. 4].
@@ Appendix B Proof of Lemma 1
Note that (2) depends on the marginal distribution of $(X,\mathbf{Y}_{k})$, (3) depends on the marginal distribution of $X$, and (5) and (6) depend on the marginal distribution of $(X,\mathbf{Z}_{l})$. Therefore, without loss of generality, using [54, Th. 3.5.2] and (52), it suffices to consider
@@ C-A Achievability Proof
Lemma 4 (Weinstein–Aronszajn Identity [52, Appx. B]).
For any $a\in\mathbb{R}_{+}$ and matrix $\mathbf{H}\in\mathbb{R}^{\Omega\times 1}$, we have
${\det}(\mathbf{H}a\mathbf{H}^{\intercal}+{\bf I}_{\Omega})=a\mathbf{H}^{\intercal}\mathbf{H}+1,$ (58)
@@
$f(\mathbf{H}^{\intercal}\mathbf{H})=\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}\mathbf{H}+1}{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}\mathbf{H}+1}\right)$ (59)
@@
Lemma 4 is applied in the calculation of mutual information with vector random variables for given $k$ and $l$, and the role of Lemma 5 is to find the minimum and maximum values of the mutual information among all possible $k\in\mathcal{K}$ and $l\in\mathcal{L}$.
@@
$\mathbf{Y}_{k}$ $=\mathbf{H}_{k}U+\mathbf{H}_{k}\Theta+\mathbf{N}_{\mathbf{Y}_{k}},$ (62)
$\mathbf{Z}_{l}$ $=\tilde{\mathbf{H}}_{l}U+\tilde{\mathbf{H}}_{l}\Theta+\mathbf{N}_{\mathbf{Z}_{l}}.$ (63)
@@
$I(\mathbf{Y}_{k};U)$ $=\frac{1}{2}\log\frac{\sigma^{2}_{X}\mathbf{H}_{k}^{\intercal}\mathbf{H}_{k}+1}{\alpha\sigma^{2}_{X}\mathbf{H}_{k}^{\intercal}\mathbf{H}_{k}+1},$ (64)
$I(\mathbf{Z}_{l};U)$ $=\frac{1}{2}\log\frac{\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1}{\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1}$ (65)
@@ C-B Converse Proof
We will need the following lemmas. These lemmas convert vector observations in (17) into scalar Gaussian random variables using sufficient statistics [46, Sect. 2.9]. This transformation plays an important role in deriving the outer bound of Gaussian sources.
Lemma 6 ([56, Lemma 3.1]).
[…]
@@
As $(X,\bar{Y}_{k},\bar{Z}_{l})$ are scalar Gaussian random variables, when the squared value of the correlation coefficient of $(X,\bar{Y}_{k^{*}})$ is greater than that of $(X,\bar{Z}_{l^{*}})$, i.e., $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, implying that $\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}\geq\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}$ for any pair $(k,l)$, there exist Gaussian random variables $(X^{\prime},\bar{Y}^{\prime}_{k},\bar{Z}^{\prime}_{l})$ such that $X^{\prime}-\bar{Y}^{\prime}_{k}-\bar{Z}^{\prime}_{l}$ is satisfied, and the marginal distributions of $(X,\bar{Y}_{k})$ and $(X^{\prime},\bar{Y}^{\prime}_{k})$ and that of $(X,\bar{Z}_{l})$ and $(X^{\prime},\bar{Z}^{\prime}_{l})$ coincide [47, Lemma 6]. In the subsequent discussions, we denote $(X^{\prime},\bar{Y}^{\prime}_{k},\bar{Z}^{\prime}_{l})$ as $(X,\bar{Y}_{k},\bar{Z}_{l})$ for brevity. This property is used in the derivation of the following lemma.
@@
If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, the outer bounds of the GS and CS models are provided as
$\mathcal{R}_{\rm G}$~ $\subseteq~\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bar{\mathcal{O}}_{G_{kl}},~~~\mathcal{R}_{\rm C}\subseteq~\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bar{\mathcal{O}}_{C_{kl}},$ (73)
@@
and $U$ satisfies the Markov chain $U-X-\bar{Y}_{k}-\bar{Z}_{l}$. If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}<\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, $\mathcal{R}_{\rm G}=\mathcal{R}_{\rm C}=\{(R_{S},R_{J},R_{L}):R_{S}=0,R_{J}\geq 0,R_{L}\geq 0\}.$
@@
The proof is provided in Appendix D and follows standard converse proof techniques, where Fano’s inequality and the introduction of auxiliary random variable are used. The key idea is to exploit the relationship in (70), which shows that $(\bar{Y}_{k},\bar{Z}_{l})$ and $(\mathbf{Y}_{k},\mathbf{Z}_{l})$ are mutually deterministic. This allows the vector random variables $(\mathbf{Y}_{k},\mathbf{Z}_{l})$ to be removed from the rate constraints during the analysis. ∎
@@
The conditional Fisher information of $A$ is defined by $\mathbb{J}(A|U)=\mathbb{E}\left[\left(\frac{\partial\log f_{A|U}(a|u)}{\partial a}\right)^{2}\right],$ where the expectation is taken over $(U,A)$ [49, Def. 1].
Lemma 9 ([49, Cor. 1]).
Let $W,A,B$ be random variables, and let the density for any combination of them exist. Moreover, assume that given $W$, AA and $B$ are independent. Then, we have
We use Lemma 9 to establish a lower bound on the conditional Fisher information, as presented in Lemma 10. This lemma is then used to derive a lower bound on the difference $h(\bar{Y}_{k}|U)-h(X|U)$ given that $h(\bar{Z}_{l}|U)-h(\bar{Y}_{k}|U)$ is fixed.
@@
For $0\leq r\leq 1/\nu_{\bar{Y}_{k}}$, it holds that
@@
From [49, Lemma 3], it follows that
$h(\frac{1}{\nu_{\bar{Z}_{l}}}\bar{Z}_{l}|U)-h(\frac{1}{\nu_{\bar{Y}_{k}}}\bar{Y}_{k}|U)=\frac{1}{2}\int_{1/\nu_{\bar{Y}_{k}}}^{1/\nu_{\bar{Z}_{l}}}\mathbb{J}(X+\sqrt{t}\tilde{N}|U)dt$ (84)
@@
$\frac{1}{2}$ $\int_{1/\nu_{\bar{Y}_{k}}}^{1/\nu_{\bar{Z}_{l}}}\mathbb{J}(X+\sqrt{t}\tilde{N}|U)dt$
$\overset{\rm(a)}{=}\frac{1}{2}\int_{1/\nu_{\bar{Y}_{k}}}^{1/\nu_{\bar{Z}_{l}}}\mathbb{J}(X+\sqrt{r}N+\sqrt{t-r}N^{\prime}|U)dt$
@@
$=\frac{1}{2}\int_{1/\nu_{\bar{Y}_{k}}}^{1/\nu_{\bar{Z}_{l}}}\frac{\mathbb{J}(X+\sqrt{r}N|U)}{1+\mathbb{J}(X+\sqrt{r}N|U)(t-r)}dt$
@@
where (a) follows by picking a real number $r$ in the range of $0\leq r\leq 1/\nu_{\bar{Y}_{k}}$ and using independent Gaussian random variables $N\sim\mathcal{N}(0,1)$ and $N^{\prime}\sim\mathcal{N}(0,1)$, and (b) is due to Lemma 9. Lastly, comparing (80) and (85), we obtain (83). ∎
@@
$\mathcal{R}_{\rm G}$ $\subseteq\bigcup_{0<\alpha\leq 1}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$~
@@ Appendix D Proof of Lemma 8
where the left-hand sides of (90) and (91), and (92) hold because $\bar{Y}^{n}_{k}$ and $\bar{Z}^{n}_{l}$ are functions of $\mathbf{Y}^{n}_{k}$ and $\mathbf{Z}^{n}_{l}$, respectively, by Lemma 7, and the right-hand sides of (90) and (91) are due to the sufficient statistic [46, Sect. 2.9]. In addition, for the scalar random variables $(X,\bar{Y}_{k},\bar{Z}_{l})$, the Markov chain $X^{n}-\bar{Y}^{n}_{k}-\bar{Z}^{n}_{l}$ holds for any pair $(k,l)$ when $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$. Combining this with (92) gives
@@
$\overset{\rm(b)}{=}H(S|J,\mathbf{Z}^{n}_{l},\bar{Z}^{n}_{l})-H(S|J,\mathbf{Y}^{n}_{k},\bar{Y}^{n}_{k})+n(\delta+\delta_{n})$
$\overset{\rm(c)}{=}H(S|J,\bar{Z}^{n}_{l})-H(S|J,\bar{Y}^{n}_{k})+n(\delta+\delta_{n})$
@@
$\overset{\rm(d)}{=}\sum_{t=1}^{n}\{I(\bar{Y}_{k,t};U_{t}|V_{t})-I(\bar{Z}_{l,t};U_{t}|V_{t})\}+n(\delta+\delta_{n})$ (97)
$\overset{\rm(e)}{=}\sum_{t=1}^{n}\{I(\bar{Y}_{k,t};U_{t})-I(\bar{Z}_{l,t};U_{t})$
@@
where (a) is due to (5) and Fano’s inequality with $\delta_{n}$ defined in (96) as the secret key $S$ can be reliably estimated from $(J,\mathbf{Y}^{n}_{k})$, (b) follows from the left-hand sides of (90) and (91), (c) holds by the right-hand sides of (90) and (91), (d) follows by [7, Lemma 4.1], (e) follows from the Markov chains $V_{t}-U_{t}-\bar{Y}_{k,t}$ and $V_{t}-U_{t}-\bar{Z}_{l,t}$, and (f) is due to (95), which results in $I(\bar{Y}_{k,t};V_{t})-I(\bar{Z}_{l,t};V_{t})\geq 0$.
@@
$\overset{\rm(c)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k})-n\delta_{n}$
$\overset{\rm(d)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k})-n\delta_{n}$ (99)
$\overset{\rm(e)}{=}\sum_{t=1}^{n}\{h(X_{t}|\bar{Y}_{k,t})-h(X_{t}|J,S,X^{t-1},\bar{Y}^{n}_{k},\bar{Z}^{t-1}_{l})\}-n\delta_{n}$
@@
$I(X^{n};J|\mathbf{Z}^{n}_{l})$ $\overset{\rm(a)}{=}I(X^{n};J)-I(\mathbf{Z}^{n}_{l};J)$
$\overset{\rm(b)}{=}I(X^{n};J)-I(\mathbf{Z}^{n}_{l},\bar{Z}^{n}_{l};J)$
$\overset{\rm(c)}{=}I(X^{n};J)-I(\bar{Z}^{n}_{l};J)$
$\overset{\rm(d)}{=}I(X^{n};J|\bar{Z}^{n}_{l}),$ (101)
@@
$\overset{\rm(b)}{=}I(X^{n};J,S,\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-n\delta_{n}$
$\overset{\rm(c)}{=}I(X^{n};J,S,\bar{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\bar{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-n\delta_{n}$
$\overset{\rm(d)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k},\bar{Z}^{n}_{l})+I(X^{n};\bar{Y}^{n}_{k}|\bar{Z}^{n}_{l})$
@@
where (a) is due to (101), (b) follows from (70), i.e., $\bar{Y}^{n}_{k}$ is a function of $\mathbf{Y}^{n}_{k}$, (c) holds because the Markov chain $X^{n}-(J,S,\bar{Y}^{n}_{k},\bar{Z}^{n}_{l})-\mathbf{Y}^{n}_{k}$ holds, from the right-hand side of (90), (d) and (e) are due to the Markov chain $(J,S)-X^{n}-\bar{Y}^{n}_{k}-\bar{Z}^{n}_{l}$, obtained from (93), and (e) follows by the same steps from (99) to (100).